Circular No. 5989

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994P IN UGC 6983 Penny Sackett, Institute for Advanced Studies, Princeton University,

reports her discovery of an apparent supernova on spectrograms of UGC 6983 taken by A. Gould (Ohio State University) with the 1.8-m telescope at Lowell Observatory on Apr. 20. The long-slit spectrum was centered on Hα, with the long slit aligned with the major axis of the galaxy. In addition to the spectrum of the galaxy, there appears a spatially unresolved feature with very strong, broad emission around Hα, asymmetric and having width of ∼ 3000 km/s; this unresolved feature is much brighter than the Hα emission of the galaxy. SN 1994P appears ∼ 120′′ northeast of the galaxy’s center (toward p.a. 75^{o}). A. Filippenko, University of California at Berkeley, writes: “P. Sackett faxed me an uncalibrated spectrum of her supernova candidate in UGC 6983. The object does indeed seem to be a type-II supernova, several months past maximum brightness. Hα is purely (or nearly purely) in emission.” M. W. Richmond, Princeton University, measured the following position for SN 1994P from a V -band image taken by T. Matheson (University of California, Berkeley) with the Nickel 1-m reflector at Lick Observatory on May 14 UT: α = 11^{h}56^{m}46^{s}.57 ± 0^{s}.04, δ = +52^{o}59′40′′.7 ± 0′′.1 (equinox 1950.0); the corresponding offset from the galaxy ‘nucleus’ (defined as the bright patch in the central bar of the galaxy, very slightly east of the geometric center) is 105′′.2 east and 31′′.1 north. Richmond also measured V ∼ 18.2 for the supernova.

SUPERNOVA 1994I IN NGC 5194 L. V. Morrison and R. W. Argyle, Royal Greenwich Observatory, re-

port an accurate optical position of SN 1994I obtained with the Carlsberg Automatic Meridian Circle on La Palma (operated by L. Helmer and R. Argyle): α = 13^{h}29^{m}54^{s}.072, δ = +47^{o}11′30′′.50 (equinox J2000.0, FK5 reference frame; epoch 1994.28; ± 0′′.08 in both coordinates; from six observations in α and nine in δ). They add: “This optical position is ∼ 0′′.5 west and 0′′.1 north of the VLA position reported on IAUC 5963; this optical position should not be biased towards the nucleus of the galaxy, which is ∼ 14′′ west and 12′′ north of the supernova. However, a gradient in the background intensity around SN 1994I could bias our result. This would produce a systematic shift of the position as the magnitude of the supernova declined. There is no evidence of this in the individual observations.”

1994 May 17 Daniel W. E. Green